# Work statistics for Quantum Spin Chains: characterizing quantum phase transitions, benchmarking time evolution, and examining passivity of quantum states

### v3

@@ Abstract
We study three aspects of work statistics in the context of the fluctuation theorem for the quantum spin chains by numerical methods based on matrix-product states. First, we elaborate that the work done on the spin-chain by a sudden quench can be used to characterize the quantum phase transitions (QPT). We further obtain the numerical results to demonstrate its capability of characterizing the QPT of both Landau-Ginzbrug types, such as the Ising chain, or topological types, such as the Haldane chain. Second, we propose to use the fluctuation theorem, such as Jarzynski’s equality, which relates the real-time correlator to the ratio of the thermal partition functions, as a benchmark indicator for the numerical real-time evolving methods. Third, we study the passivity of ground and thermal states of quantum spin chains under some cyclic impulse processes. We show that the passivity of thermal states and ground states under the hermitian actions are ensured by the second laws and variational principles, respectively, and also verify it by numerical calculations. Besides, we also consider the passivity of ground states under non-hermitian actions, for which the variational principle cannot be applied. Despite that, we find no violation of passivity from our numerical results for all the cases considered in both Ising-like and Haldane-like chains.
@@ I Introduction
Motivated by the progress of quantum simulation in the cold atom experiments [1, 2, 3, 4, 5] and the theoretical understanding of the thermalization of isolated quantum systems [6, 7, 8], the nonequilibrium dynamics of quantum systems have been studied extensively [9, 10, 11, 12, 13]. A system can be driven into nonequilibrium simply by introducing a time-dependent interaction or a time-dependent coupling constant, and the system can return to equilibrium after turning off the time dependence of the coupling. For a many-body system, obtaining the exact dynamical evolution is generally difficult. Given a numerical method of evaluating the dynamical evolution, finding a way of estimating the numerical accuracy becomes a challenging task, especially for a system without prior knowledge of the exact dynamics. One would expect that the numerical error accumulates as the system evolves. Therefore, a reliable real-time error estimator will help compare the numerical results with the experimental ones, such as the quantum simulation. On the other hand, for a sudden quenching process, one can bypass the above difficulty.
@@
For nonequilibrium dynamics of an isolated quantum system, one would expect that the underlying microreversibility should manifest in some way, in contrast with the macroscopic second law of thermodynamics. Indeed, equality relations exist, collectively known as fluctuation theorem [14, 15, 16, 17, 18, 19, 20, 21, 22, 23], for such manifestation. The work done in a nonequilibrium process is not a state variable and will depend on the path connecting the initial and final state. The fluctuation theorem relates the average work to the free energy difference between the initial and final states. To calculate the average work or the higher moments, one can construct the corresponding characteristic function (or generating function of work statistics), which can be rewritten as the real-time correlation function [17, 19, 24]. This characteristic function of work takes a form similar to the out-of-time-ordered correlator (OTOC) in characterizing the quantum chaos [25, 26, 27]. In general, one needs to adopt some numerical method to calculate this generating function or some OTOC-like quantities involving nonequilibrium dynamical evaluation, which may be spoiled by the accumulation of numerical errors. On the other hand, the difference in free energy only depends on the initial and final states. It will require disproportionately less numerical effort than the one needed for the characteristic function. Exploiting the feature of such disproportionality in the numerical efforts on both sides of the fluctuation theorem, one can use it to monitor the accuracy of any numerical method for evaluating the real-time nonequilibrium dynamical evolution of a quantum system. In recent works [28, 29], a numerical method based on the matrix-product-state (MPS) approach [30, 31] is adopted to evaluate the characteristic function for spin chains and verify the fluctuation theorem. This method can also be generalized to 2-dimensional spin-lattices, called the tensor-network state (TNS) [32]. In this paper, we will adopt the MPS method to evaluate the characteristic function of other quantum spin chains and use the fluctuation theorem to benchmark the capability of these numerical time-evolving methods.
@@
Another interesting aspect of work statistics is characterizing the (dynamical) phase transitions [33, 34, 33]. For such purpose, one can consider the sudden quench, which is realized by a sudden change of the Hamiltonian. In this case, it is straightforward to see that the average work done on the isolated system measures the expectation value of the jump of the Hamiltonian across the quenching point with respect to the initial state. This can be seen as measuring the excitation energy of the initial state induced by the sudden change of the Hamiltonian. When considering the quantum phase transition, the initial state is the ground state. After the sudden quench, it becomes an excited state to the new Hamiltonian. If both initial and final Hamiltonians are non-critical and gapped, then only the gapped excited states with their energies less than the average work done on the system can be induced. However, if the final Hamiltonian is the one at the quantum critical point, it describes a gapless system, then an infinite number of gapless excitations will be induced. We expect the work statistics to differ from those for the non-critical cases because this large number of gapless excitations will provide more dynamic paths for the quenching process for retrieving a definite amount of work. This implies that the average work excitation can display discontinuous behavior when the sudden quench connects a non-critical Hamiltonian to a critical one. In this sense, the work statistics, i.e., the average work done or the higher cumulants, can be adopted to characterize the quantum phase transitions, i.e., as an order parameter. Since the above reasoning concerns only the gapless feature of the critical systems, it can work for both the quantum phase transition of the Landau-Ginzburg type due to spontaneously-symmetry-breaking (SSB) or the topological type characterized by nonlocal order parameters. Finding the appropriate order parameter to indicate the topological phase transition is usually more difficult. In this paper, we will demonstrate that work done by sudden quench can be used to identify both types of quantum phase transitions.
@@
Based on the work statistics and the fluctuation theorem, we will use reliable numerical methods to investigate the above three aspects of the quantum spin lattice models in this paper: characterizing the quantum phase transitions, benchmarking the accuracy of the numerical real-time evolution, and examining the passivity. These studies enlarge the perspective of the work done and the fluctuation theorem.
@@ II.1 Work statistics and fluctuation theorem
Motivated by the work statistics as the characteristic of thermal or quantum fluctuations in the context of fluctuation theorem. We first have a brief sketch of the basics.
@@
The denominator is unity for unitary $U(t_{f})$ (or hermitian $H(t)$). The expression (10) can be recast into the real-time correlation function on the extended Schwinger-Keldysh contour [24] so that techniques of open system dynamics can be adopted. We usually omit $t_{f}$ and write $G(u)$.
@@
where we choose the initial Hamiltonian $H_{0}=H(0^{-})$ to be time-independent and hermitian; however, the driving operator $V$ is time-independent but could be non-hermitian. If $V$ is non-hermitian, then $U(t)$ is non-unitary so that the denominator of (10) is not unity. An initial state $\rho_{0}$, which is taken to be either the thermal state or ground state of $H_{0}$, will be driven away from equilibrium due to the nontrivial time-dependence of the coupling constant $\lambda(t)$. The notations $0^{-}$ and $0^{+}$ used later denote the moments right before and after $t=0$, respectively.
@@ II.2 Benchmark the numerical real-time evolution
Jarzynski’s equality (9) is an interesting relationship that relates a real-time correlator to the ratio of the partition functions, i.e., $e^{-\beta\Delta F}={Z(t_{f})\over Z(0)}$. There is usually no analytical method to calculate the real-time correlators such as $G(u)$ of Eq. (10) of a many-body system even though there might be some analytical ways of calculating the partition function. Even relying on the numerical method to evaluate the real-time correlators, the numerical errors accumulate more as the evolution continues. On the other hand, the partition function is a stationary quantity free of the accumulated error. Therefore, we can characterize the accumulated error by defining the following ratio:
@@ II.3 Work done by sudden quench and phase transition
In general, the work statistics or its characteristic function for the many-body system is difficult to evaluate for the complication of real-time dynamics. However, to characterize the phase transition, we can bypass the difficulty by just considering the work done by a sudden quench implemented by the Hamiltonian of (11) and (12). Denote the ground state of $H_{0}$ by $\rho_{0}\equiv|0\rangle\langle 0|$, and the Hmailtoian after quench by $H_{+}$, i.e., $H_{+}\equiv H_{0}+\Delta\lambda V$, then the characteristic function can be simplified as follows:
@@
which yields
$\overline{W}$ $=$ ${\rm Tr}[\rho_{0}\Delta H]=\Delta\lambda\langle 0|V|0\rangle\;,$ (16)
$\sigma^{2}_{W}$ $=$ $\langle 0|H_{+}^{2}|0\rangle-\langle 0|H_{+}|0\rangle^{2}\;,$
$=$ $(\Delta\lambda)^{2}\Big{(}\langle 0|V^{2}|0\rangle-\langle 0|V|0\rangle^{2}\Big{)}\;.$
The quantity $\overline{W}$, as shown for the sudden quench, measures the energy of the ”excited state” $|0\rangle$ with respect to $H_{+}$, the quantity $\sigma_{W}$ measures the corresponding fluctuation of $\Delta H=\Delta\lambda V$. Thus, $\overline{W}$ (or $\sigma^{2}_{W}$) could act as a non-local order parameter to characterize the quantum phase transition while tuning $H_{0}$.
We now elaborate on why the quantities in Eq. (16) and Eq. (II.3) can be used as the order parameter to characterize the phase transition. First, if the initial and final states are Gibbs states, then the Hamiltonian can also be understood as the modular Hamiltonian, i.e., $H_{\rm mod}=-\log\rho$. In this case, $\overline{W}$ can be rewritten as the relative entropy $S_{\rm rel}$, i.e.,
@@
The phase transition is usually characterized by the order parameter, which is the expectation value of some local or non-local observables. We can decompose the Hamiltonian $H_{+}=\sum_{f}E_{f^{+}}|f^{+}\rangle\langle f^{+}|$, so that (16) can be rewritten as
@@
The average work done $\overline{W}$ can be seen as the expectation value of $E_{b^{+}}$ with its probability measure given by $p(f)=|\langle 0|f^{+}\rangle|^{2}$. We will tune some coupling in $H_{0}$ over a range covering different phases. Again, $|0\rangle$ looks like the excited state to $H_{+}$. When $H_{0}$ and $H_{+}$ are in the same phase, we expect $|0^{+}\rangle\approx|0\rangle$ so that $p(f)=\delta_{f,0}$ so that $\overline{W}\simeq E_{0^{+}}|\langle 0|0^{+}\rangle|^{2}\simeq E_{0^{+}}$. On the other hand, if they are in different phases, we should not expect a sharp distribution for $p(f)$, and ${\overline{W}}$ will be quite different from $E_{0^{+}}$ due to the incoherent average. Moreover, when $H_{+}$ is the gapless critical Hamiltonian at the quantum critical point, we expect the emergence of a dense set of degenerate gapless states denoted as $\{0^{+}_{c}\}$ such that $p(f)$ has sizeable values within this set. Thus, $\overline{W}\simeq\sum_{\{0^{+}_{c}\}}E_{0^{+}}p(c)$, which gives a sudden jump from the coherent value of $E_{0^{+}}$ to the one of incoherent. Therefore, $\overline{W}$ can be used as an order parameter for phase transition. Similar arguments go for $\sigma^{2}_{W}$ or higher moments.
@@ II.4 Passivity of a quantum state: thermal or ground states
Therefore, the fluctuation theorem guarantees the thermal states’ passivity, thus, the second law. Since the second law is established only in the thermodynamics limit, there is evidence that the thermal-like states of finite systems under some peculiar process can be active [41, 42, 43].
Intuitively, the passivity of the ground states should hold because the operation $V$ excites the ground state to the states with larger energies so that one cannot extract energy from the ground state. We can check passivity by first obtaining the average work extraction $W_{\rm ext}^{\rm cyclic}\equiv-\overline{W}$ for the cyclic process with $H(t_{f})=H(0):=H_{0}$ from the characteristic function of (10), i.e.,
@@
Note that ${\rm Tr}\rho_{1}={\rm Tr}\rho_{2}=1$ Moreover, if $U(t_{f})$ is unitary, i.e., $U^{\dagger}(t_{f})=U^{-1}(t_{f})$, then $\rho_{1}=\rho_{0}$ and $\rho_{2}=U(t_{f})\rho_{0}U^{\dagger}(t_{f})$, so that (21) can be reduced to
@@
which is nothing but the energy difference between the ground state $\rho_{0}$ and the excited state $U(t_{f})\rho_{0}U^{\dagger}(t_{f})$. Thus, in such cases, the passivity of the ground state is guaranteed if the variational principle holds. On the other hand, if $U(t_{f})$ is non-unitary, then $\rho_{1}\neq\rho_{0}$ even though $\rho_{2}$ bears the look of excited state under non-unitary evolution. Therefore, the negativity of $W_{\rm ext}^{\rm cyclic}$ of (21) can no longer be ensured by the variational principle.
We can conclude from the above discussions that the variational principle generally ensures the ground state’s passivity. In this work, we will verify this by the numerical calculation and explore its possible violations by considering the cases of non-hermitian $V$ so that the variational principle cannot be applied to guarantee the negativity of $W_{\rm ext}^{\rm cyclic}$ of (21).
For simplicity and numerical implementation, we will only consider the impulse cyclic processes, i.e., with $\lambda(t)$ given by (13), for which we have
@@
It turns out that we can easily implement MPS formalism for spin chain models to evaluate the right side of (26) numerically, and then extract $W_{\rm ext}^{\rm impulse}$ from the results. Alternatively, the average work extraction can also be obtained from (21)-(23) by replacing $U(t_{f})$ with $e^{-i\lambda V}$ and $U^{\dagger}(t_{f})$ with $e^{i\lambda V^{\dagger}}$.

### v4

@@ Abstract
We study three aspects of work statistics in the context of the fluctuation theorem for the quantum spin chains up to $1024$ sites by numerical methods based on matrix-product states (MPS). First, we use our numerical method to evaluate the moments/cumulants of work done by sudden quench process on the Ising or Haldane spin chains and study their behaviors across the quantum phase transitions. Our results show that, up to the fourth cumulant, the work statistics can indicate the quantum phase transition characterized by the local order parameters but barely for purely topological phase transitions. Second, we propose to use the fluctuation theorem, such as Jarzynski’s equality, which relates the real-time correlator to the ratio of the thermal partition functions, as a benchmark indicator for the numerical real-time evolving methods. Third, we study the passivity of ground and thermal states of quantum spin chains under some cyclic impulse processes. We show that the passivity of thermal states and ground states under the hermitian actions are ensured by the second laws and variational principles, respectively, and also verify it by numerical calculations. Besides, we also consider the passivity of ground states under non-hermitian actions, for which the variational principle cannot be applied. Despite that, we find no violation of passivity from our numerical results for all the cases considered in the Ising and Haldane chains. Overall, we demonstrate that the work statistics for the sudden quench and impulse processes can be evaluated precisely by the numerical MPS method to characterize quantum phase transitions and examine the passivity of quantum states. We also propose to exploit the universality of the fluctuation theorem to benchmark the numerical real-time evolutions in an algorithm and model independent way.
@@ I Introduction
Motivated by the progress of quantum simulation in the cold atom experiments [1, 2, 3, 4, 5] and the theoretical understanding of the thermalization of isolated quantum systems [6, 7, 8], the non-equilibrium dynamics of quantum systems have been studied extensively [9, 10, 11, 12, 13]. A system can be driven into nonequilibrium simply by introducing a time-dependent interaction or a time-dependent coupling constant, and the system can return to equilibrium after turning off the time dependence of the coupling. For a many-body system, obtaining the exact dynamical evolution is generally difficult. Given a numerical method of evaluating the dynamical evolution, finding a way of estimating the numerical accuracy becomes a challenging task, especially for a system without prior knowledge of the exact dynamics. One would expect that the numerical error accumulates as the system evolves. Therefore, a reliable real-time error estimator will help compare the numerical results with the experimental ones, such as the quantum simulation.
@@
For nonequilibrium dynamics of an isolated quantum system, one would expect that the underlying microreversibility should manifest in some way, in contrast with the macroscopic second law of thermodynamics. Indeed, equality relations exist, collectively known as fluctuation theorem [14, 15, 16, 17, 18, 19, 20, 21, 22, 23], for such manifestation. The work done in a nonequilibrium process is not a state variable and will depend on the path connecting the initial and final state. The fluctuation theorem relates the average work to the free energy difference between the initial and final states. To calculate the average work or the higher moments, one can construct the corresponding characteristic function (or generating function of work statistics), which can be rewritten as the real-time correlation function [17, 19, 24]. This characteristic function of work takes a form similar to the out-of-time-ordered correlator (OTOC) in characterizing the quantum chaos [25, 26, 27]. In general, one needs to adopt some numerical method to calculate this generating function or some OTOC-like quantities involving nonequilibrium dynamical evaluation, which may be spoiled by the accumulation of numerical errors. On the other hand, the difference in free energy only depends on the initial and final states. It will require disproportionately less numerical effort than the one needed for the characteristic function. Exploiting the feature of such disproportionality in the numerical efforts on both sides of the fluctuation theorem, one can use it to monitor the accuracy of any numerical method for evaluating the real-time nonequilibrium dynamical evolution of a quantum system. In recent works [28], a numerical method based on the matrix-product-state (MPS) approach [29, 30] is adopted to evaluate the characteristic function for Ising spin chains and verify the fluctuation theorem, see also [31] for using the similar method for evaluating the heat statistics of open systems. This method can also be generalized to 2-dimensional spin-lattices, called the tensor-network state (TNS) [32]. In this paper, we will adopt the MPS method to evaluate the characteristic function of other quantum spin chains and use the fluctuation theorem to benchmark the capability of these numerical time-evolving methods.
Another interesting aspect of work statistics is characterizing the (dynamical) quantum phase transitions (QPTs). This was firstly demonstrated in [33] for the Ising chain and later in [34] for the random matrix models under a sudden quench process, i.e, with a sudden change of the Hamiltonian, for which the work characteristic function is related to the Loschmidt echo. Moreover, the average work done on the isolated system measures the expectation value of the jump of the Hamiltonian across the quenching point with respect to the initial state. Then, a similar trick is applied to show the universal scaling behavior of the kink statistics for the dynamical quantum phase transition under the Kibble-Zurek mechanism (KZM) for the (Ising) spin chains [35, 36, 37, 38], and for Kitave honeycomb model [39]. This approach to characterizing QPTs by work statistics is extended to the (exactly solvable) spin/fermion chain models under the non-quench processes [35, 40, 41].
It is also known that entanglement structures like entanglement entropy can also characterize QPTs by showing a discontinuity at critical points in many cases of QPTs, e.g., see [42, 43]. As both the entanglement entropy and the work statistics can be encoded as some combinations of multi-point correlations, this may explain why both can characterize QPTs through some singular multi-point correlations. This perspective has been explored in [44] for the fermionic Hubbard model with random impurities. It showed that the entanglement is minimized while the work average is maximized at the critical point, but the second moment of work vanishes. In this paper, we will study this issue further for the Ising and Haldane chains.
@@
When considering the quantum phase transition, the initial state is the ground state. After the sudden quench, it becomes a linear combination of the various excited states of the new Hamiltonian close in energy. If both initial and final Hamiltonians are non-critical and gapped, then only the gapped excited states with their energies less than the average work done on the system can be induced. However, if the final Hamiltonian happens to be at the quantum critical point, it describes a gapless system, then an infinite number of gapless excitations will be induced. We expect the work statistics to differ from those for the non-critical cases because this large number of gapless excitations will provide more dynamic paths for the quenching process for retrieving a definite amount of work. This implies that the average work excitation can display discontinuous behavior when the sudden quench connects a non-critical Hamiltonian to a critical one. In this sense, the work statistics, i.e., the average work done or the higher cumulants, can be adopted to characterize the quantum phase transitions, i.e., as an order parameter. Since the above reasoning concerns only the gapless feature of the critical systems, it can work for both the quantum phase transition of the Landau-Ginzburg type due to spontaneously-symmetry-breaking (SSB) or the topological type characterized by nonlocal order parameters. Finding the appropriate order parameter to indicate the topological phase transition is usually more difficult. In this paper, we will explore the power of the numerical MPS-based algorithm to evaluate the moments/cumulants of work statistics and demonstrate their capability and limitations in characterizing the QPTs of the quantum Ising and Haldane chains. With the help of the quantum-state RG of MPS, we can evaluate the work statistics for the chains up to $1024$ sites. This will effectively suppress the finite-size effect, especially when near the quantum critical point.
@@
Based on the work statistics and the fluctuation theorem, we will use reliable numerical methods to investigate the above three aspects of the quantum spin lattice models in this paper: characterizing the quantum phase transitions, benchmarking the accuracy of the numerical real-time evolution, and examining the passivity. These studies enlarge the perspective of the work done and the fluctuation theorem. In summary, in this work we will demonstrate the power of the numerical MPS method in evaluating the work statistics precisely enough to characterize the quantum phase transitions and examine the passivity of quantum states, even for the non-Hermitian actions. Moreover, we will also show how to exploit the simplicity and universality of the fluctuation theorem to benchmark the numerical real-time evolutions for generic models and numerical algorithms.
@@ II.1 Work statistics and fluctuation theorem
We first give a brief sketch of the basics of work statistics in the context of the fluctuation theorem, which provides an elegant framework for describing and characterizing thermal or quantum fluctuations in statistical mechanics.
@@
The denominator is unity for unitary $U(t_{f})$ (or hermitian $H(t)$). The expression of Eq. (10) can be recast into the real-time correlation function on the extended Schwinger-Keldysh contour [24] so that techniques of open system dynamics can be adopted. We usually omit $t_{f}$ and write $G(u)$.
@@
where we choose the initial Hamiltonian $H_{0}=H(0^{-})$ to be time-independent and hermitian; however, the driving operator $V$ is time-independent but could be non-hermitian. If $V$ is non-hermitian, then $U(t)$ is non-unitary so that the denominator of Eq. (10) is not unity. An initial state $\rho_{0}$, which is taken to be either the thermal state or ground state of $H_{0}$, will be driven away from equilibrium due to the nontrivial time-dependence of the coupling constant $\lambda(t)$. The notations $0^{-}$ and $0^{+}$ used later denote the moments right before and after $t=0$, respectively.
@@ II.2 Benchmark the numerical real-time evolution
Jarzynski’s equality shown in Eq. (9) is an interesting relationship that relates a real-time correlator to the ratio of the partition functions, i.e., $e^{-\beta\Delta F}={Z(t_{f})\over Z(0)}$. There is usually no analytical method to calculate the real-time correlators such as $G(u)$ of Eq. (10) of a many-body system even though there might be some analytical ways of calculating the partition function. Even relying on the numerical method to evaluate the real-time correlators, the numerical errors accumulate more as the evolution continues. On the other hand, the partition function is a stationary quantity free of the accumulated error. Therefore, we can characterize the accumulated error by defining the following ratio:
@@ II.3 Work done by sudden quench and phase transition
In general, the work statistics or its characteristic function for the many-body system is difficult to evaluate for the complication of real-time dynamics. However, to characterize the phase transition, we can bypass the difficulty by just considering the work done by a sudden quench implemented by the Hamiltonian of Eq. (11) and Eq. (12). Denote the ground state of $H_{0}$ by $\rho_{0}\equiv|0\rangle\langle 0|$, and the Hmailtoian after quench by $H_{+}$, i.e., $H_{+}\equiv H_{0}+\Delta\lambda V$, then the characteristic function can be simplified as follows:
@@
which yields moments and the first few cumulants of the work done as follows,
$\overline{W^{m}}$ $=$ $(\Delta\lambda)^{m}\;\langle 0|V^{m}|0\rangle\;,$ (16)
$\sigma^{2}_{W}$ $=$ $\langle 0|H_{+}^{2}|0\rangle-\langle 0|H_{+}|0\rangle^{2}\;,$ (17)
$=$ $(\Delta\lambda)^{2}\Big{(}\langle 0|V^{2}|0\rangle-\langle 0|V|0\rangle^{2}\Big{)}\;,$
$\kappa_{3}$ $=$ $\overline{W^{3}}-3\overline{W^{2}}\overline{W}+\overline{W}^{3}\;,$ (18)
$\kappa_{4}$ $=$ $\overline{W^{4}}-4\overline{W^{3}}\overline{W}-3\overline{W^{2}}^{2}+12\overline{W^{2}}\overline{W}^{2}-6\overline{W}^{4}\;.\qquad$ (19)
The quantity $\overline{W}$, as shown for the sudden quench, measures the energy of the ”excited state” $|0\rangle$ with respect to $H_{+}$, the quantity $\sigma_{W}$ measures the corresponding fluctuation of $\Delta H=\Delta\lambda V$. Thus, $\overline{W}$ (or $\sigma^{2}_{W}$) could act as a local order parameter to characterize the quantum phase transition while tuning $H_{0}$. From Eq. (16) we can also calculate the higher moments $\overline{W^{m}}$ or higher cumulants $\kappa_{m}$ for the sudden quench in terms of the expectation value of higher order $V^{m}$’s.
We now elaborate on why the quantities in Eq. (16) and Eq. (17) can be used as the order parameter to characterize the phase transition. First, suppose the initial and final states are Gibbs states related by unitary evolution. In that case, the Hamiltonian can also be understood as the modular Hamiltonian, i.e., $H_{\rm mod}=-\log\rho$. In this case, $\overline{W}$ can be rewritten as the relative entropy $S_{\rm rel}$ [20], i.e.,
@@
The phase transition is usually characterized by the order parameter, which is the expectation value of some local or non-local observables. We can decompose the Hamiltonian $H_{+}=\sum_{f}E_{f^{+}}|f^{+}\rangle\langle f^{+}|$, so that Eq. (16) with $m=1$ can be rewritten as
@@
The average work done $\overline{W}$ can be seen as the expectation value of $E_{f^{+}}$ with its probability measure given by $p(f)=|\langle 0|f^{+}\rangle|^{2}$. We will tune some coupling in $H_{0}$ over a range covering different phases. Again, $|0\rangle$ looks like the excited state to $H_{+}$. When $H_{0}$ and $H_{+}$ are in the same phase, we expect $|0^{+}\rangle\approx|0\rangle$ so that $p(f)=\delta_{f,0}$ so that $\overline{W}\simeq E_{0^{+}}|\langle 0|0^{+}\rangle|^{2}\simeq E_{0^{+}}$. On the other hand, if they are in different phases, we should not expect a sharp distribution for $p(f)$, and ${\overline{W}}$ will be quite different from $E_{0^{+}}$ due to the incoherent average. For example, when $H_{+}$ is the gapless critical Hamiltonian at the quantum critical point, we expect the emergence of a dense set of degenerate gapless states denoted as $\{0^{+}_{c}\}$ such that the density of ground states $p(0^{+})=\sum_{\{0^{+}_{c}\}}p(c)<1$, which is quite different from the case of non-critical gapped $H_{+}$ with $p(0^{+})\simeq 1$. This then induces a sudden jump of $\overline{W}$ from $\overline{W}\simeq E_{0}^{+}$ to the incoherent average $\overline{W}\simeq E_{0^{+}}\sum_{\{0^{+}_{c}\}}p(0^{+}_{c})+\sum_{f\neq\{0^{+}\}}p(f)E_{f}$. Therefore, $\overline{W}$ can be used as an order parameter for phase transition. Similar arguments go for $\sigma^{2}_{W}$ or higher moments/cumulants. Some preliminary study of $p(f)$ for quantum Ising chain under sudden quench by the exact diagonalization (ED) method is done in Appendix A, and the result is shown in Fig. 11 therein.
@@ II.4 Passivity of a quantum state: thermal or ground states
Therefore, the fluctuation theorem guarantees the thermal states’ passivity, thus, the second law. Since the second law is established only in the thermodynamic limit, there is evidence that the thermal-like states of finite systems under some peculiar process can be active [51, 52, 53].
Based on the above result, it is worthy of mentioning the study in [54], which considered the passivity of the energy eigenstates of Ising chains under the action of local Hermitian operations. By adopting the eigenstate thermalization hypothesis [55, 56, 8], an energy eigenstate locally looks like a thermal state, e.g., the reduced density matrix of a local region is approximately a thermal one with some effective temperature. Thus, based on the passivity of the thermal state and ETH, an energy eigenstate will probably be passive under the action of local unitary operations. In [54], it was shown that all the energy eigenstates are passive under local operations for non-integrable Ising chains of finite size, while most energy eigenstates are passive for integrable Ising chains. This is consistent with the fact that ETH holds only weakly for integrable systems, see, e.g., [57, 58, 59]. The above result could also help to explain the passivity of the ground states studied below and imply a negative effective temperature.
Intuitively, the ground states should be passive because the operation $V$ excites the ground state to those with larger energies so that one cannot extract energy from the ground state. We can check passivity by first obtaining the average work extraction $W_{\rm ext}^{\rm cyclic}\equiv-\overline{W}$ for the cyclic process with $H(t_{f})=H(0):=H_{0}$ from the characteristic function of Eq. (10), i.e.,
@@
Note that ${\rm Tr}\rho_{1}={\rm Tr}\rho_{2}=1$ Moreover, if $U(t_{f})$ is unitary, i.e., $U^{\dagger}(t_{f})=U^{-1}(t_{f})$, then $\rho_{1}=\rho_{0}$ and $\rho_{2}=U(t_{f})\rho_{0}U^{\dagger}(t_{f})$, so that Eq. (23) can be reduced to
@@
which is nothing but the energy difference between the ground state $\rho_{0}$ and the excited state $U(t_{f})\rho_{0}U^{\dagger}(t_{f})$. Thus, in such cases, the passivity of the ground state is guaranteed if the variational principle holds. On the other hand, if $U(t_{f})$ is non-unitary, then $\rho_{1}\neq\rho_{0}$ even though $\rho_{2}$ bears the look of excited state under non-unitary evolution. Therefore, the negativity of $W_{\rm ext}^{\rm cyclic}$ of Eq. (23) can no longer be ensured by the variational principle.
We can conclude from the above discussions that the variational principle generally ensures the ground state’s passivity. In this work, we will verify this by the numerical calculation and explore its possible violations by considering the cases of non-hermitian $V$ so that the variational principle cannot be applied to guarantee the negativity of $W_{\rm ext}^{\rm cyclic}$ of Eq. (23).
For simplicity and numerical implementation, we will only consider the impulse cyclic processes, i.e., with $\lambda(t)$ given by Eq. (13), for which we have
@@
It turns out that we can easily implement MPS formalism for spin chain models to evaluate the right side of Eq. (28) numerically and then extract $W_{\rm ext}^{\rm impulse}$ from the results. Alternatively, the average work extraction can also be obtained from Eq. (23)-Eq. (25) by replacing $U(t_{f})$ with $e^{-i\lambda V}$ and $U^{\dagger}(t_{f})$ with $e^{i\lambda V^{\dagger}}$.
@@
Finally, we comment on the possible physical realization of a non-Hermitian operation, although our consideration in this work is somewhat academic. There are ways to realize non-Hermitian actions. One way is to drop the quantum jump term in the Lindblad master equation for the open quantum system by choosing the proper Lindblad operator, and the resultant equation describes the Heisenberg evolution of the density matrix operator by a non-Hermitian Hamiltonian, see, e.g., [60]. This can be alternatively realized on quantum circuits by probability quantum computing with suitably dilated Hilbert space, see [61]. Or, one can introduce the non-Hermitian interaction in the context of $\mathcal{PT}$-symmetry. For the Ising chain, it is equivalent to introducing an imaginary magnetic field; see, e.g., [62].